\begin{thebibliography}{44}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Aggarwal et~al.(2023)Aggarwal, Madaan, Yang, and {Mausam}]{aggarwal2023lets}
Pranjal Aggarwal, Aman Madaan, Yiming Yang, and {Mausam}.
\newblock Let{'}s sample step by step: Adaptive-consistency for efficient reasoning and coding with {LLM}s.
\newblock In Houda Bouamor, Juan Pino, and Kalika Bali, editors, \emph{Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing}, pages 12375--12396, Singapore, December 2023. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/2023.emnlp-main.761}.
\newblock URL \url{https://aclanthology.org/2023.emnlp-main.761}.

\bibitem[AI@Meta(2024)]{llama3modelcard}
AI@Meta.
\newblock Llama 3 model card.
\newblock 2024.
\newblock URL \url{https://github.com/meta-llama/llama3/blob/main/MODEL_CARD.md}.

\bibitem[Ali et~al.(2024)Ali, Fromm, Thellmann, Rutmann, Lübbering, Leveling, Klug, Ebert, Doll, Buschhoff, Jain, Weber, Jurkschat, Abdelwahab, John, Suarez, Ostendorff, Weinbach, Sifa, Kesselheim, and Flores-Herr]{ali2024tokenizer}
Mehdi Ali, Michael Fromm, Klaudia Thellmann, Richard Rutmann, Max Lübbering, Johannes Leveling, Katrin Klug, Jan Ebert, Niclas Doll, Jasper~Schulze Buschhoff, Charvi Jain, Alexander~Arno Weber, Lena Jurkschat, Hammam Abdelwahab, Chelsea John, Pedro~Ortiz Suarez, Malte Ostendorff, Samuel Weinbach, Rafet Sifa, Stefan Kesselheim, and Nicolas Flores-Herr.
\newblock Tokenizer choice for llm training: Negligible or crucial?, 2024.

\bibitem[Bisk et~al.(2020)Bisk, Zellers, Bras, Gao, and Choi]{bisk2020piqa}
Yonatan Bisk, Rowan Zellers, Ronan~Le Bras, Jianfeng Gao, and Yejin Choi.
\newblock {PIQA:} reasoning about physical commonsense in natural language.
\newblock In \emph{The Thirty-Fourth {AAAI} Conference on Artificial Intelligence, {AAAI} 2020, The Thirty-Second Innovative Applications of Artificial Intelligence Conference, {IAAI} 2020, The Tenth {AAAI} Symposium on Educational Advances in Artificial Intelligence, {EAAI} 2020, New York, NY, USA, February 7-12, 2020}, pages 7432--7439. {AAAI} Press, 2020.
\newblock \doi{10.1609/AAAI.V34I05.6239}.
\newblock URL \url{https://doi.org/10.1609/aaai.v34i05.6239}.

\bibitem[Brown et~al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal, Neelakantan, Shyam, Sastry, Askell, Agarwal, Herbert-Voss, Krueger, Henighan, Child, Ramesh, Ziegler, Wu, Winter, Hesse, Chen, Sigler, Litwin, Gray, Chess, Clark, Berner, McCandlish, Radford, Sutskever, and Amodei]{brown2020language}
Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared~D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel Ziegler, Jeffrey Wu, Clemens Winter, Chris Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei.
\newblock Language models are few-shot learners.
\newblock In H.~Larochelle, M.~Ranzato, R.~Hadsell, M.F. Balcan, and H.~Lin, editors, \emph{Advances in Neural Information Processing Systems}, volume~33, pages 1877--1901. Curran Associates, Inc., 2020.
\newblock URL \url{https://proceedings.neurips.cc/paper_files/paper/2020/file/1457c0d6bfcb4967418bfb8ac142f64a-Paper.pdf}.

\bibitem[Cobbe et~al.(2021)Cobbe, Kosaraju, Bavarian, Chen, Jun, Kaiser, Plappert, Tworek, Hilton, Nakano, Hesse, and Schulman]{cobbe2021training}
Karl Cobbe, Vineet Kosaraju, Mohammad Bavarian, Mark Chen, Heewoo Jun, Lukasz Kaiser, Matthias Plappert, Jerry Tworek, Jacob Hilton, Reiichiro Nakano, Christopher Hesse, and John Schulman.
\newblock Training verifiers to solve math word problems, 2021.

\bibitem[Dagan et~al.(2024)Dagan, Synnaeve, and Rozière]{dagan2024getting}
Gautier Dagan, Gabriel Synnaeve, and Baptiste Rozière.
\newblock Getting the most out of your tokenizer for pre-training and domain adaptation, 2024.

\bibitem[Gage(1994)]{gage1994}
Philip Gage.
\newblock A new algorithm for data compression.
\newblock \emph{C Users J.}, 12\penalty0 (2):\penalty0 23–38, feb 1994.
\newblock ISSN 0898-9788.

\bibitem[{Gemma Team} et~al.(2024){Gemma Team}, Mesnard, Hardin, Dadashi, Bhupatiraju, Pathak, Sifre, Rivière, Kale, Love, Tafti, Hussenot, Sessa, Chowdhery, Roberts, Barua, Botev, Castro-Ros, Slone, Héliou, Tacchetti, Bulanova, Paterson, Tsai, Shahriari, Lan, Choquette-Choo, Crepy, Cer, Ippolito, Reid, Buchatskaya, Ni, Noland, Yan, Tucker, Muraru, Rozhdestvenskiy, Michalewski, Tenney, Grishchenko, Austin, Keeling, Labanowski, Lespiau, Stanway, Brennan, Chen, Ferret, Chiu, Mao-Jones, Lee, Yu, Millican, Sjoesund, Lee, Dixon, Reid, Mikuła, Wirth, Sharman, Chinaev, Thain, Bachem, Chang, Wahltinez, Bailey, Michel, Yotov, Chaabouni, Comanescu, Jana, Anil, McIlroy, Liu, Mullins, Smith, Borgeaud, Girgin, Douglas, Pandya, Shakeri, De, Klimenko, Hennigan, Feinberg, Stokowiec, hui Chen, Ahmed, Gong, Warkentin, Peran, Giang, Farabet, Vinyals, Dean, Kavukcuoglu, Hassabis, Ghahramani, Eck, Barral, Pereira, Collins, Joulin, Fiedel, Senter, Andreev, and Kenealy]{gemmateam2024gemma}
{Gemma Team}, Thomas Mesnard, Cassidy Hardin, Robert Dadashi, Surya Bhupatiraju, Shreya Pathak, Laurent Sifre, Morgane Rivière, Mihir~Sanjay Kale, Juliette Love, Pouya Tafti, Léonard Hussenot, Pier~Giuseppe Sessa, Aakanksha Chowdhery, Adam Roberts, Aditya Barua, Alex Botev, Alex Castro-Ros, Ambrose Slone, Amélie Héliou, Andrea Tacchetti, Anna Bulanova, Antonia Paterson, Beth Tsai, Bobak Shahriari, Charline~Le Lan, Christopher~A. Choquette-Choo, Clément Crepy, Daniel Cer, Daphne Ippolito, David Reid, Elena Buchatskaya, Eric Ni, Eric Noland, Geng Yan, George Tucker, George-Christian Muraru, Grigory Rozhdestvenskiy, Henryk Michalewski, Ian Tenney, Ivan Grishchenko, Jacob Austin, James Keeling, Jane Labanowski, Jean-Baptiste Lespiau, Jeff Stanway, Jenny Brennan, Jeremy Chen, Johan Ferret, Justin Chiu, Justin Mao-Jones, Katherine Lee, Kathy Yu, Katie Millican, Lars~Lowe Sjoesund, Lisa Lee, Lucas Dixon, Machel Reid, Maciej Mikuła, Mateo Wirth, Michael Sharman, Nikolai Chinaev, Nithum Thain, Olivier Bachem,
  Oscar Chang, Oscar Wahltinez, Paige Bailey, Paul Michel, Petko Yotov, Rahma Chaabouni, Ramona Comanescu, Reena Jana, Rohan Anil, Ross McIlroy, Ruibo Liu, Ryan Mullins, Samuel~L Smith, Sebastian Borgeaud, Sertan Girgin, Sholto Douglas, Shree Pandya, Siamak Shakeri, Soham De, Ted Klimenko, Tom Hennigan, Vlad Feinberg, Wojciech Stokowiec, Yu~hui Chen, Zafarali Ahmed, Zhitao Gong, Tris Warkentin, Ludovic Peran, Minh Giang, Clément Farabet, Oriol Vinyals, Jeff Dean, Koray Kavukcuoglu, Demis Hassabis, Zoubin Ghahramani, Douglas Eck, Joelle Barral, Fernando Pereira, Eli Collins, Armand Joulin, Noah Fiedel, Evan Senter, Alek Andreev, and Kathleen Kenealy.
\newblock Gemma: Open models based on gemini research and technology, 2024.

\bibitem[Gorenflo and Crano(1998)]{gorenflo1998multiple}
Daniel~W Gorenflo and William~D Crano.
\newblock The multiple perspectives inventory: A measure of perspective-taking.
\newblock \emph{Swiss Journal of Psychology}, 57\penalty0 (3):\penalty0 163--177, 1998.

\bibitem[Groeneveld et~al.(2024)Groeneveld, Beltagy, Walsh, Bhagia, Kinney, Tafjord, Jha, Ivison, Magnusson, Wang, Arora, Atkinson, Authur, Chandu, Cohan, Dumas, Elazar, Gu, Hessel, Khot, Merrill, Morrison, Muennighoff, Naik, Nam, Peters, Pyatkin, Ravichander, Schwenk, Shah, Smith, Strubell, Subramani, Wortsman, Dasigi, Lambert, Richardson, Zettlemoyer, Dodge, Lo, Soldaini, Smith, and Hajishirzi]{groeneveld2024olmo}
Dirk Groeneveld, Iz~Beltagy, Pete Walsh, Akshita Bhagia, Rodney Kinney, Oyvind Tafjord, Ananya~Harsh Jha, Hamish Ivison, Ian Magnusson, Yizhong Wang, Shane Arora, David Atkinson, Russell Authur, Khyathi~Raghavi Chandu, Arman Cohan, Jennifer Dumas, Yanai Elazar, Yuling Gu, Jack Hessel, Tushar Khot, William Merrill, Jacob Morrison, Niklas Muennighoff, Aakanksha Naik, Crystal Nam, Matthew~E. Peters, Valentina Pyatkin, Abhilasha Ravichander, Dustin Schwenk, Saurabh Shah, Will Smith, Emma Strubell, Nishant Subramani, Mitchell Wortsman, Pradeep Dasigi, Nathan Lambert, Kyle Richardson, Luke Zettlemoyer, Jesse Dodge, Kyle Lo, Luca Soldaini, Noah~A. Smith, and Hannaneh Hajishirzi.
\newblock Olmo: Accelerating the science of language models, 2024.

\bibitem[Gu and Dao(2023)]{gu2023mamba}
Albert Gu and Tri Dao.
\newblock Mamba: Linear-time sequence modeling with selective state spaces, 2023.

\bibitem[Guo et~al.(2019)Guo, Zheng, Fan, Yu, and Wang]{hao2019visual}
Hao Guo, Kang Zheng, Xiaochuan Fan, Hongkai Yu, and Song Wang.
\newblock Visual attention consistency under image transforms for multi-label image classification.
\newblock In \emph{2019 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR)}, pages 729--739, 2019.
\newblock \doi{10.1109/CVPR.2019.00082}.

\bibitem[Han et~al.(2018)Han, Yao, Yu, Niu, Xu, Hu, Tsang, and Sugiyama]{bo2018coteaching}
Bo~Han, Quanming Yao, Xingrui Yu, Gang Niu, Miao Xu, Weihua Hu, Ivor Tsang, and Masashi Sugiyama.
\newblock Co-teaching: Robust training of deep neural networks with extremely noisy labels.
\newblock In S.~Bengio, H.~Wallach, H.~Larochelle, K.~Grauman, N.~Cesa-Bianchi, and R.~Garnett, editors, \emph{Advances in Neural Information Processing Systems}, volume~31. Curran Associates, Inc., 2018.
\newblock URL \url{https://proceedings.neurips.cc/paper_files/paper/2018/file/a19744e268754fb0148b017647355b7b-Paper.pdf}.

\bibitem[Hendrycks et~al.(2021)Hendrycks, Burns, Kadavath, Arora, Basart, Tang, Song, and Steinhardt]{hendrycksmath2021}
Dan Hendrycks, Collin Burns, Saurav Kadavath, Akul Arora, Steven Basart, Eric Tang, Dawn Song, and Jacob Steinhardt.
\newblock Measuring mathematical problem solving with the math dataset.
\newblock \emph{NeurIPS}, 2021.

\bibitem[Hewitt et~al.(2022)Hewitt, Manning, and Liang]{hewitt2022truncation}
John Hewitt, Christopher Manning, and Percy Liang.
\newblock Truncation sampling as language model desmoothing.
\newblock In Yoav Goldberg, Zornitsa Kozareva, and Yue Zhang, editors, \emph{Findings of the Association for Computational Linguistics: EMNLP 2022}, pages 3414--3427, Abu Dhabi, United Arab Emirates, December 2022. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/2022.findings-emnlp.249}.
\newblock URL \url{https://aclanthology.org/2022.findings-emnlp.249}.

\bibitem[Holtzman et~al.(2020)Holtzman, Buys, Du, Forbes, and Choi]{Holtzman2020The}
Ari Holtzman, Jan Buys, Li~Du, Maxwell Forbes, and Yejin Choi.
\newblock The curious case of neural text degeneration.
\newblock In \emph{International Conference on Learning Representations}, 2020.
\newblock URL \url{https://openreview.net/forum?id=rygGQyrFvH}.

\bibitem[Jain et~al.(2023)Jain, Schwarzschild, Wen, Somepalli, Kirchenbauer, yeh Chiang, Goldblum, Saha, Geiping, and Goldstein]{jain2023baseline}
Neel Jain, Avi Schwarzschild, Yuxin Wen, Gowthami Somepalli, John Kirchenbauer, Ping yeh Chiang, Micah Goldblum, Aniruddha Saha, Jonas Geiping, and Tom Goldstein.
\newblock Baseline defenses for adversarial attacks against aligned language models, 2023.

\bibitem[Jain et~al.(2024)Jain, Ma, Deoras, and Xiang]{jain2024lightweight}
Siddhartha Jain, Xiaofei Ma, Anoop Deoras, and Bing Xiang.
\newblock Lightweight reranking for language model generations, 2024.

\bibitem[Jiang et~al.(2023)Jiang, Sablayrolles, Mensch, Bamford, Chaplot, de~las Casas, Bressand, Lengyel, Lample, Saulnier, Lavaud, Lachaux, Stock, Scao, Lavril, Wang, Lacroix, and Sayed]{jiang2023mistral}
Albert~Q. Jiang, Alexandre Sablayrolles, Arthur Mensch, Chris Bamford, Devendra~Singh Chaplot, Diego de~las Casas, Florian Bressand, Gianna Lengyel, Guillaume Lample, Lucile Saulnier, Lélio~Renard Lavaud, Marie-Anne Lachaux, Pierre Stock, Teven~Le Scao, Thibaut Lavril, Thomas Wang, Timothée Lacroix, and William~El Sayed.
\newblock Mistral 7b, 2023.

\bibitem[Kojima et~al.(2022)Kojima, Gu, Reid, Matsuo, and Iwasawa]{kojima2022large}
Takeshi Kojima, Shixiang~Shane Gu, Machel Reid, Yutaka Matsuo, and Yusuke Iwasawa.
\newblock Large language models are zero-shot reasoners.
\newblock In Alice~H. Oh, Alekh Agarwal, Danielle Belgrave, and Kyunghyun Cho, editors, \emph{Advances in Neural Information Processing Systems}, 2022.
\newblock URL \url{https://openreview.net/forum?id=e2TBb5y0yFf}.

\bibitem[Kudo(2018)]{kudo2018subword}
Taku Kudo.
\newblock Subword regularization: Improving neural network translation models with multiple subword candidates.
\newblock In Iryna Gurevych and Yusuke Miyao, editors, \emph{Proceedings of the 56th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, pages 66--75, Melbourne, Australia, July 2018. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/P18-1007}.
\newblock URL \url{https://aclanthology.org/P18-1007}.

\bibitem[Li et~al.(2023)Li, Lin, Zhang, Fu, Chen, Lou, and Chen]{li2023making}
Yifei Li, Zeqi Lin, Shizhuo Zhang, Qiang Fu, Bei Chen, Jian-Guang Lou, and Weizhu Chen.
\newblock Making language models better reasoners with step-aware verifier.
\newblock In Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki, editors, \emph{Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, pages 5315--5333, Toronto, Canada, July 2023. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/2023.acl-long.291}.
\newblock URL \url{https://aclanthology.org/2023.acl-long.291}.

\bibitem[Li et~al.(2024{\natexlab{a}})Li, Yuan, Feng, Pan, Sun, Wang, Wang, and Li]{li2024turnig}
Yiwei Li, Peiwen Yuan, Shaoxiong Feng, Boyuan Pan, Bin Sun, Xinglin Wang, Heda Wang, and Kan Li.
\newblock Turning dust into gold: Distilling complex reasoning capabilities from llms by leveraging negative data.
\newblock In Michael~J. Wooldridge, Jennifer~G. Dy, and Sriraam Natarajan, editors, \emph{Thirty-Eighth {AAAI} Conference on Artificial Intelligence, {AAAI} 2024, Thirty-Sixth Conference on Innovative Applications of Artificial Intelligence, {IAAI} 2024, Fourteenth Symposium on Educational Advances in Artificial Intelligence, {EAAI} 2014, February 20-27, 2024, Vancouver, Canada}, pages 18591--18599. {AAAI} Press, 2024{\natexlab{a}}.
\newblock \doi{10.1609/AAAI.V38I17.29821}.
\newblock URL \url{https://doi.org/10.1609/aaai.v38i17.29821}.

\bibitem[Li et~al.(2024{\natexlab{b}})Li, Yuan, Feng, Pan, Wang, Sun, Wang, and Li]{li2024escape}
Yiwei Li, Peiwen Yuan, Shaoxiong Feng, Boyuan Pan, Xinglin Wang, Bin Sun, Heda Wang, and Kan Li.
\newblock Escape sky-high cost: Early-stopping self-consistency for multi-step reasoning.
\newblock In \emph{The Twelfth International Conference on Learning Representations}, 2024{\natexlab{b}}.
\newblock URL \url{https://openreview.net/forum?id=ndR8Ytrzhh}.

\bibitem[Ling et~al.(2017)Ling, Yogatama, Dyer, and Blunsom]{ling2017program}
Wang Ling, Dani Yogatama, Chris Dyer, and Phil Blunsom.
\newblock Program induction by rationale generation: Learning to solve and explain algebraic word problems.
\newblock In Regina Barzilay and Min-Yen Kan, editors, \emph{Proceedings of the 55th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, pages 158--167, Vancouver, Canada, July 2017. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/P17-1015}.
\newblock URL \url{https://aclanthology.org/P17-1015}.

\bibitem[Liu et~al.(2023)Liu, Deng, Sabour, Jia, Huang, and Mihalcea]{liu2023task}
Siyang Liu, Naihao Deng, Sahand Sabour, Yilin Jia, Minlie Huang, and Rada Mihalcea.
\newblock Task-adaptive tokenization: Enhancing long-form text generation efficacy in mental health and beyond.
\newblock In Houda Bouamor, Juan Pino, and Kalika Bali, editors, \emph{Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing}, pages 15264--15281, Singapore, December 2023. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/2023.emnlp-main.944}.
\newblock URL \url{https://aclanthology.org/2023.emnlp-main.944}.

\bibitem[Meister et~al.(2023)Meister, Pimentel, Wiher, and Cotterell]{meister2023locally}
Clara Meister, Tiago Pimentel, Gian Wiher, and Ryan Cotterell.
\newblock Locally typical sampling.
\newblock \emph{Transactions of the Association for Computational Linguistics}, 11:\penalty0 102--121, 2023.
\newblock \doi{10.1162/tacl_a_00536}.
\newblock URL \url{https://aclanthology.org/2023.tacl-1.7}.

\bibitem[Nagata(1994)]{nagata1994stochastic}
Masaaki Nagata.
\newblock A stochastic {J}apanese morphological analyzer using a forward-{DP} backward-{A}* n-best search algorithm.
\newblock In \emph{{COLING} 1994 Volume 1: The 15th {I}nternational {C}onference on {C}omputational {L}inguistics}, Kyoto, Japan, August 1994.
\newblock URL \url{https://aclanthology.org/C94-1032}.

\bibitem[Naik et~al.(2024)Naik, Chandrasekaran, Yuksekgonul, Palangi, and Nushi]{naik2024diversity}
Ranjita Naik, Varun Chandrasekaran, Mert Yuksekgonul, Hamid Palangi, and Besmira Nushi.
\newblock Diversity of thought improves reasoning abilities of llms, 2024.

\bibitem[Penedo et~al.(2024)Penedo, Kydlíček, von Werra, and Wolf]{penedo2024fineweb}
Guilherme Penedo, Hynek Kydlíček, Leandro von Werra, and Thomas Wolf.
\newblock Fineweb, April 2024.
\newblock URL \url{https://huggingface.co/datasets/HuggingFaceFW/fineweb}.

\bibitem[Provilkov et~al.(2020)Provilkov, Emelianenko, and Voita]{provilkov2020bpe}
Ivan Provilkov, Dmitrii Emelianenko, and Elena Voita.
\newblock {BPE}-dropout: Simple and effective subword regularization.
\newblock In Dan Jurafsky, Joyce Chai, Natalie Schluter, and Joel Tetreault, editors, \emph{Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics}, pages 1882--1892, Online, July 2020. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/2020.acl-main.170}.
\newblock URL \url{https://aclanthology.org/2020.acl-main.170}.

\bibitem[Runco and Acar(2012)]{runco2012divergent}
Mark~A Runco and Selcuk Acar.
\newblock Divergent thinking as an indicator of creative potential.
\newblock \emph{Creativity research journal}, 24\penalty0 (1):\penalty0 66--75, 2012.

\bibitem[Scott(2002)]{scott2002bayesian}
Steven~L. Scott.
\newblock Bayesian methods for hidden markov models: Recursive computing in the 21st century.
\newblock \emph{Journal of the American Statistical Association}, 97\penalty0 (457):\penalty0 337--351, 2002.
\newblock ISSN 01621459.
\newblock URL \url{http://www.jstor.org/stable/3085787}.

\bibitem[Sennrich et~al.(2016)Sennrich, Haddow, and Birch]{sennrich2016neural}
Rico Sennrich, Barry Haddow, and Alexandra Birch.
\newblock Neural machine translation of rare words with subword units.
\newblock In Katrin Erk and Noah~A. Smith, editors, \emph{Proceedings of the 54th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, pages 1715--1725, Berlin, Germany, August 2016. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/P16-1162}.
\newblock URL \url{https://aclanthology.org/P16-1162}.

\bibitem[Singh and Strouse(2024)]{singh2024tokenization}
Aaditya~K. Singh and DJ~Strouse.
\newblock Tokenization counts: the impact of tokenization on arithmetic in frontier llms, 2024.

\bibitem[Soldaini et~al.(2024)Soldaini, Kinney, Bhagia, Schwenk, Atkinson, Authur, Bogin, Chandu, Dumas, Elazar, Hofmann, Jha, Kumar, Lucy, Lyu, Lambert, Magnusson, Morrison, Muennighoff, Naik, Nam, Peters, Ravichander, Richardson, Shen, Strubell, Subramani, Tafjord, Walsh, Zettlemoyer, Smith, Hajishirzi, Beltagy, Groeneveld, Dodge, and Lo]{dolma}
Luca Soldaini, Rodney Kinney, Akshita Bhagia, Dustin Schwenk, David Atkinson, Russell Authur, Ben Bogin, Khyathi Chandu, Jennifer Dumas, Yanai Elazar, Valentin Hofmann, Ananya~Harsh Jha, Sachin Kumar, Li~Lucy, Xinxi Lyu, Nathan Lambert, Ian Magnusson, Jacob Morrison, Niklas Muennighoff, Aakanksha Naik, Crystal Nam, Matthew~E. Peters, Abhilasha Ravichander, Kyle Richardson, Zejiang Shen, Emma Strubell, Nishant Subramani, Oyvind Tafjord, Pete Walsh, Luke Zettlemoyer, Noah~A. Smith, Hannaneh Hajishirzi, Iz~Beltagy, Dirk Groeneveld, Jesse Dodge, and Kyle Lo.
\newblock {Dolma: an Open Corpus of Three Trillion Tokens for Language Model Pretraining Research}.
\newblock \emph{arXiv preprint}, 2024.

\bibitem[Stickland and Murray(2020)]{stickland2020diverse}
Asa~Cooper Stickland and Iain Murray.
\newblock Diverse ensembles improve calibration, 2020.

\bibitem[Wang et~al.(2023)Wang, Wei, Schuurmans, Le, Chi, Narang, Chowdhery, and Zhou]{wang2023selfconsistency}
Xuezhi Wang, Jason Wei, Dale Schuurmans, Quoc~V Le, Ed~H. Chi, Sharan Narang, Aakanksha Chowdhery, and Denny Zhou.
\newblock Self-consistency improves chain of thought reasoning in language models.
\newblock In \emph{The Eleventh International Conference on Learning Representations}, 2023.
\newblock URL \url{https://openreview.net/forum?id=1PL1NIMMrw}.

\bibitem[Wang et~al.(2006)Wang, Dogan, and Lin]{wang2007mp}
Yan Wang, Enis Dogan, and Xiaodong Lin.
\newblock The effect of multiple-perspective thinking on problem solving.
\newblock In \emph{Proceedings of the 7th International Conference on Learning Sciences}, ICLS '06, page 812–817. International Society of the Learning Sciences, 2006.
\newblock ISBN 0805861742.

\bibitem[Wei et~al.(2022)Wei, Wang, Schuurmans, Bosma, brian ichter, Xia, Chi, Le, and Zhou]{wei2022chain}
Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, brian ichter, Fei Xia, Ed~H. Chi, Quoc~V Le, and Denny Zhou.
\newblock Chain of thought prompting elicits reasoning in large language models.
\newblock In Alice~H. Oh, Alekh Agarwal, Danielle Belgrave, and Kyunghyun Cho, editors, \emph{Advances in Neural Information Processing Systems}, 2022.
\newblock URL \url{https://openreview.net/forum?id=_VjQlMeSB_J}.

\bibitem[Yao et~al.(2023)Yao, Yu, Zhao, Shafran, Griffiths, Cao, and Narasimhan]{yao2023tree}
Shunyu Yao, Dian Yu, Jeffrey Zhao, Izhak Shafran, Thomas~L. Griffiths, Yuan Cao, and Karthik Narasimhan.
\newblock Tree of thoughts: Deliberate problem solving with large language models, 2023.

\bibitem[Yeo et~al.(2021)Yeo, Kar, and Zamir]{yeo2021robustness}
Teresa Yeo, O\u{g}uzhan~Fatih Kar, and Amir Zamir.
\newblock Robustness via cross-domain ensembles.
\newblock In \emph{Proceedings of the IEEE/CVF International Conference on Computer Vision (ICCV)}, pages 12189--12199, October 2021.

\bibitem[Zouhar et~al.(2023)Zouhar, Meister, Gastaldi, Du, Vieira, Sachan, and Cotterell]{zouhar2023formal}
Vilém Zouhar, Clara Meister, Juan~Luis Gastaldi, Li~Du, Tim Vieira, Mrinmaya Sachan, and Ryan Cotterell.
\newblock A formal perspective on byte-pair encoding, 2023.

\end{thebibliography}

\documentclass{article}

\pdfoutput=1

\usepackage{microtype}
\usepackage{graphicx}
\usepackage{subfigure}
\usepackage{booktabs} 

\usepackage{hyperref}

\usepackage[accepted]{icml2023}

\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{mathtools}
\usepackage{amsthm}

\usepackage[capitalize,noabbrev]{cleveref}

\theoremstyle{plain}

\theoremstyle{definition}

\theoremstyle{remark}

\usepackage[textsize=tiny]{todonotes}

\usepackage{amsmath}
\usepackage{graphicx}
\usepackage{multicol}
\usepackage{multirow}
\usepackage{makecell}
\usepackage{csquotes}
\usepackage{url}

\newcommand{\olm}{\textsc{OLMo-7B}}
\newcommand{\gmtwo}{\textsc{Gemma-2B}}
\newcommand{\gmsev}{\textsc{Gemma-7B}}
\newcommand{\lmate}{\textsc{LLaMA3-8B}}
\newcommand{\lmsev}{\textsc{LLaMA3-70B}}
\newcommand{\mis}{\textsc{Mistral-7B}}
\newcommand{\mam}{\textsc{Mamba-2.8B}}
\newcommand{\seq}[2]{\{#1^1,\dots,#1^{#2}\}}
\newcommand{\seqt}[2][1]{[t^{#1},\dots,t^{#2}]}
\newcommand{\vocab}{\mathcal{V}}
\newcommand{\xstr}{\mathbf{X}}
\newcommand{\xtok}{\mathbf{x}_\text{tok}}
\newcommand{\xalltok}{\mathbf{T}(\xstr)}
\newcommand{\dataset}{\mathcal{D}}
\newcommand{\trainset}{\dataset_\text{train}}

\newcommand{\ystr}{\mathbf{Y}}

\definecolor{navyblue}{HTML}{091F92}
\hypersetup{
    colorlinks=true,
    allcolors=navyblue,
}

\icmltitlerunning{Improving Self Consistency in LLMs through Probabilistic Tokenization}

\begin{document}

\twocolumn[
\icmltitle{Improving Self Consistency in LLMs through Probabilistic Tokenization}



\icmlsetsymbol{equal}{*}

\begin{icmlauthorlist}
\icmlauthor{Ashutosh Sathe}{equal,yyy,comp}
\icmlauthor{Divyanshu Aggarwal}{equal,comp}
\icmlauthor{Sunayana Sitaram}{comp}
\icmlaffiliation{yyy}{IIT Bombay}
\icmlaffiliation{comp}{Microsoft Research India}
\end{icmlauthorlist}


\icmlcorrespondingauthor{Ashutosh Sathe}{absathe@cse.iitb.ac.in}

\icmlkeywords{Machine Learning, ICML}

\vskip 0.3in
]



\printAffiliationsAndNotice{\icmlEqualContribution} 

\begin{abstract}
Prior research has demonstrated noticeable performance gains through the use of probabilistic tokenizations, an approach that involves employing multiple tokenizations of the same input string during the training phase of a language model. Despite these promising findings, modern large language models (LLMs) have yet to be trained using probabilistic tokenizations. Interestingly, while the tokenizers of these contemporary LLMs have the capability to generate multiple tokenizations, this property remains underutilized.

In this work, we propose a novel method to leverage the multiple tokenization capabilities of modern LLM tokenizers, aiming to enhance the self-consistency of LLMs in reasoning tasks. Our experiments indicate that when utilizing probabilistic tokenizations, LLMs generate logically diverse reasoning paths, moving beyond mere surface-level linguistic diversity. 
We carefully study probabilistic tokenization and offer insights to explain the self consistency improvements it brings through extensive experimentation on 5 LLM families and 4 reasoning benchmarks. 
\end{abstract}

\section{Introduction}

Being able to view a problem from multiple standpoints has been shown to be correlated to higher problem solving abilities and creativity \citep{gorenflo1998multiple,wang2007mp,runco2012divergent}.
In contrast, modern large language models (LLMs) such as \mis\ \citep{jiang2023mistral}, \olm\ \citep{groeneveld2024olmo}, \mam\ \citep{gu2023mamba} etc. often view language as a \textit{unique} sequence of tokens. Byte-Pair Encoding \citep{gage1994, sennrich2016neural} is a popular tokenization method employed by many LLMs to convert a given string into a sequence of tokens. Tokens in Byte-Pair Encoding (BPE) are often  ``merges'' of smaller tokens which are also present in the vocabulary. E.g. token ``\_token'' could be a merge of tokens ``\_to'' and ``ken''. This means that there are multiple valid tokenizations of a given string depending on which (sub)tokens the tokenizer encoding function chooses to merge. Table \ref{tab:intro} illustrates the phenomenon in action.

\begin{table*}[!ht] 
    \centering
    \small
    \caption{\textbf{Multiple tokenizations of a given sentence using \mis\ BPE tokenizer.} The original input string (top row) can be tokenized into multiple possible sequences of valid tokens from the \mis\ vocabulary. ``--'' represents whitespace and ``$\rvert$'' is used to indicate token boundary. The sequence of token IDs is also presented below the tokenization.}
    \begin{tabular}{p{0.95\textwidth}}
        \toprule
        A sentence can have multiple tokenizations with the BPE or Unigram tokenizer.\\ 
        \midrule        --A$\rvert$--sentence$\rvert$--can$\rvert$--have$\rvert$--multiple$\rvert$--token$\rvert$izations$\rvert$--with$\rvert$--the$\rvert$\textcolor{red}{--}$\rvert$\textcolor{red}{BP}$\rvert$\textcolor{red}{E}$\rvert$--or$\rvert$\textcolor{blue}{--Un}$\rvert$\textcolor{blue}{i}$\rvert$\textcolor{blue}{gram}$\rvert$\textcolor{magenta}{--token}$\rvert$\textcolor{magenta}{izer}$\rvert$. \\
        \{330, 12271, 541, 506, 5166, 6029, 13809, 395, 272, \textcolor{red}{28705}, \textcolor{red}{9399}, \textcolor{red}{28749}, 442, \textcolor{blue}{935}, \textcolor{blue}{28710}, \textcolor{blue}{1596}, \textcolor{magenta}{6029}, \textcolor{magenta}{4024}, 28723\}\\
        \midrule
        --A$\rvert$--sentence$\rvert$--can$\rvert$--have$\rvert$--multiple$\rvert$--token$\rvert$izations$\rvert$--with$\rvert$--the$\rvert$\textcolor{red}{--B}$\rvert$\textcolor{red}{PE}$\rvert$--or$\rvert$\textcolor{blue}{--U}$\rvert$\textcolor{blue}{ni}$\rvert$\textcolor{blue}{gram}$\rvert$\textcolor{magenta}{--token}$\rvert$\textcolor{magenta}{ize}$\rvert$\textcolor{magenta}{r}$\rvert$.\\
        \{330, 12271, 541, 506, 5166, 6029, 13809, 395, 272, \textcolor{red}{365}, \textcolor{red}{1767}, 442, \textcolor{blue}{500}, \textcolor{blue}{3023}, \textcolor{blue}{1596}, \textcolor{magenta}{6029}, \textcolor{magenta}{653}, \textcolor{magenta}{28712}, 28723\}\\
        \midrule
        --A$\rvert$--sentence$\rvert$--can$\rvert$--have$\rvert$--multiple$\rvert$--token$\rvert$izations$\rvert$--with$\rvert$--the$\rvert$\textcolor{red}{--B}$\rvert$\textcolor{red}{PE}$\rvert$--or$\rvert$\textcolor{blue}{--Un}$\rvert$\textcolor{blue}{igr}$\rvert$\textcolor{blue}{am}$\rvert$\textcolor{magenta}{--to}$\rvert$\textcolor{magenta}{ken}$\rvert$\textcolor{magenta}{izer}$\rvert$. \\
\{330, 12271, 541, 506, 5166, 6029, 13809, 395, 272, \textcolor{red}{365}, \textcolor{red}{1767}, 442, \textcolor{blue}{935}, \textcolor{blue}{3421}, \textcolor{blue}{314}, \textcolor{magenta}{298}, \textcolor{magenta}{2314}, \textcolor{magenta}{4024}, 28723\}\\
        \bottomrule
    \end{tabular}
    \label{tab:intro}
\end{table*}

Prior work \citep{kudo2018subword, provilkov2020bpe} on ``subword regularization'' has shown that including such multiple tokenizations when training neural machine translators helps the model learn better token embeddings and become robust to noisy inputs. While modern LLMs are often trained \textit{without} subword regularization, their tokenizers (mainly BPE with byte-fallback) maintain the ability to generate multiple tokenizations for a given input string as shown in Table \ref{tab:intro}. In this work, we aim to use these multiple tokenizations to improve self consistency in LLM reasoning.

An intuitive way to improve self consistency of an LLM in a reasoning task is to generate diverse reasoning paths \citep{wang2023selfconsistency} for a given problem. This is parallel to the multi-perspective reasoning as studied in \citet{wang2007mp}. Given these multiple reasoning paths, \citet{wang2023selfconsistency} propose to select the answer produced by majority of reasoning paths as the final answer. Crucially, \citet{wang2023selfconsistency} and many works that follow \citep{aggarwal2023lets,li2024turnig, jain2024lightweight, li2024escape} rely on diversity promoting text generation techniques such as Nucleus sampling \citep{Holtzman2020The} or temperature sampling to get diversity in the reasoning paths. 

In this work, we propose to use the multiple tokenizations as a primary way to generate diverse reasoning paths. Our hypothesis is that different sequences of tokens for the same input string should naturally lead to different and diverse generations. This is advantageous since unlike prior sampling methods \citep{Holtzman2020The, hewitt2022truncation, meister2023locally}, it does not rely on the model's next-token distribution to have sufficient diversity. 
It is important to note that we need a principled approach to generate multiple tokenizations of the given string since randomly dropping $p$\% of BPE merges can lead to degradation in performance as shown by \citet{jain2023baseline}. We extend \citet{kudo2018subword} and present an approach (Section \ref{sec:probtok}) to assign likelihood to a given tokenization and sampling a tokenization proportional to its likelihood. 

\section{Probabilistic Tokenization}
\label{sec:probtok}
Given an input string $\xstr$ and an existing vocabulary $\vocab$, we want to sample $m$ different tokenizations $\seq{\xtok}{m}$ such that each $\xtok^i$ is sampled proportional to $\Pr(\xtok^i|\xstr)$. Each tokenization $\xtok^i$ is a sequence of tokens $\seqt{k_i}$ where each $t^j \in \vocab$ and decoding $\xtok^i$ using $\vocab$ gives $\xstr$ back.

\subsection{Sampling a Tokenization}

We follow \citet{kudo2018subword} and use a unigram language model to estimate $\Pr(\xtok^i|\xstr)$. This means we can write $\Pr(\xtok^i|\xstr)$ using the unigram probabilities $p(t^j)$ as,

\begin{equation}
\begin{aligned}
    \Pr(\xtok^i|\xstr) &= \prod_{j=1}^{k_i}p(t^j), \\ &\text{ where } \xtok^i = \seqt{k_i}, \sum_{t^j \in \vocab}p(t^j) = 1
\end{aligned}
    \label{eq:unigram}
\end{equation}

Given the vocabulary $\vocab$ and dataset $\trainset$ of sentences, \citet{kudo2018subword} propose using the Expectation Maximization algorithm to estimate $p(t^j)$. This EM algorithm aims to maximize the marginal likelihood over the entire dataset considering $p(t^j)$ as latent variables. If $\xalltok$ denotes all possible tokenizations of a given sentence $\xstr$, the marginal can be written as,

\begin{equation}
\mathcal{L} = \sum_{s=1}^{|\trainset|}\log\Pr(\xstr_s) = \sum_{s=1}^{|\trainset|}\log\Pr\left(\sum_{\xtok\in\xalltok}\Pr(\xtok)\right)
\end{equation}

In practice, we opted for a simple counting based method to estimate $p(t^j)$. For every document in the $\trainset$, we first obtain a tokenization $\xtok^\text{BPE}$ using the existing BPE tokenizer and simply count the total occurrences of $t^j \in \xtok^\text{BPE}$ to estimate $p(t^j)$ as $\log p(t^j) = \log (\text{counts}(t^j) / N)$. Here, $N$ indicates the total number of BPE tokens produced by the tokenizer on the entire $\trainset$. For special tokens such as beginning/end of sequence, padding or unknown tokens, we set $\log p(t^j) = 0$. We provide additional discussion on counting vs EM based estimation in Appendix \ref{app:counting_vs_em}.

Once the unigram likelihoods are estimated for the entire vocabulary, we can efficiently get $l$-best tokenizations according to $\Pr(\xtok|\xstr)$ using the Forward-DP Backward-A* algorithm \citep{nagata1994stochastic}. Following \citet{kudo2018subword}, we sample from these $l$ tokenizations as $\Pr(\xtok^i|\xstr) \sim \Pr(\xtok^i)^\alpha / \sum_{i=1}^{l} \Pr(\xtok^i)^\alpha$ with $\alpha$ as a smoothing coefficient. We can now sample $m$ tokenizations for a given $\xstr$ from the best $l\geq m$-best tokenizations.

\citet{kudo2018subword} also suggests a way to accurately sample from \textit{all} ($l\rightarrow\infty$) possible tokenizations using Forward-Filtering and Backward-Sampling algorithm \citep{scott2002bayesian}. We study effects of both $l\rightarrow\infty$ and fixed $l$ on our sampled tokenizations and downstream tasks. We find that setting $l$ to a fixed value or a value dependent on $m$ can lead to less or superficial diversity in the reasoning paths sampled. More discussion on this is presented in Appendix \ref{app:finite_vs_infinite}.

\subsection{Evaluating Models with Multiple Tokenizations}\label{sec:probtok_eval_strats}

We can use probabilistic tokenization to improve self-consistency of LLMs on various types of tasks. In this work, we are interested in specifically interested in improving self consistency of reasoning based tasks.

In these tasks, we expect the model to reason through a problem (chain-of-thought) step-by-step \citep{wei2022chain} before producing the final answer. \citet{wang2023selfconsistency} has shown that sampling diverse reasoning paths from the model and picking the most common answer greatly improves the performance. While \citet{wang2023selfconsistency} rely on sampling methods such as Nucleus sampling \citep{Holtzman2020The}, we propose to use probabilistic tokenization as a way to boost diversity even further. Sampling based methods rely on the next-token distribution to be sufficiently diverse in order to generate diverse reasoning paths. As opposed to this, probabilistic tokenization directly changes the input to the model (i.e. the sequence of tokens) which should naturally lead to diverse generations. Given the reasoning problem as an input string $\xstr$, we sample $m$ different tokenizations $\seq{\xtok}{m}$ as described above and generate $m$ reasoning paths leading to $m$ different answers as $\seq{\ystr_\text{pred}}{m}$. The final answer is selected using majority vote similar to \citet{wang2023selfconsistency}. We also report ``Oracle'' task accuracy where we consider 
$\xstr$ solved if \textit{any} one of the $\seq{\ystr_\text{pred}}{m}$ matches the gold answer.

\section{Experiments}
\label{sec:experiments}


We conduct experiments to evaluate the efficacy of the proposed probabilistic tokenization on various reasoning tasks. Our findings suggest that probabilistic tokenization can robustly improve the reasoning capabilities of many LLMs.

\subsection{Setup}

\paragraph{Probabilistic Tokenization} We use a subset of the \textsc{FineWeb} dataset \citep{penedo2024fineweb} consisting of roughly 10B tokens to estimate $p(t^j)$. For both reasoning and log-likelihood based tasks, we use $l\rightarrow\infty$ i.e. we sample $m$ different from all possible tokenizations. The ablations on fixed $l$ are presented in Appendix \ref{app:finite_vs_infinite}.

\paragraph{Language Models} We experiment with four transformer-based language model families: \olm\ \citep{groeneveld2024olmo}, \gmtwo, \gmsev\ \citep{gemmateam2024gemma}, \lmate\ \citep{llama3modelcard} and \mis\ \citep{jiang2023mistral}. We also include \mam\ \citep{gu2023mamba} as a representative non-transformer based language model.

\paragraph{Reasoning Tasks} We consider four reasoning based tasks: MATH \citep{hendrycksmath2021}, AQuA \citep{ling2017program}, GSM8k \citep{cobbe2021training} and PIQA \citep{bisk2020piqa}. For each model, we report ``Baseline'' numbers which use the standard BPE tokenization. In ``CoT + SC'' baseline, we use the chain-of-thought prompting \citep{wei2022chain} with self-consistency \citep{wang2023selfconsistency} over 64 sampled reasoning paths with standard BPE tokenization. To sample diverse reasoning paths, we set the temperature $T=0.2$ with top-$k$ ($k=64$) sampling. With ``Probabilistic Tokenization'', we use the same chain-of-thought prompt. However, we sample $m=64$ different tokenizations and use greedy decoding to generate diverse reasoning paths.


\subsection{Main Results}

\begin{table*}[!ht]
    \centering
    \small
    \caption{\textbf{Results with probabilistic tokenization on chain-of-thought based reasoning tasks.} The ``Baseline'' reports the accuracy or exact-match value with the standard BPE tokenization on that task \textit{without} any chain-of-thought prompting. All other columns report changes relative to ``Baseline''. ``Probabilistic Tokenization'' (greedy decoding) outperforms chain-of-thought with self consistency (``CoT + SC'') which uses temperature based diversity promoting sampling.  }
    \begin{tabular}{llrrrrr}
        \toprule
        Task & Model & Baseline & CoT + SC & CoT + SC & Probabilistic  & Probabilistic  \\ 
        ~ & ~ & ~ & Majority & Oracle & Majority & Oracle \\
        \midrule
        \multirow{7}{*}{\bf MATH} & \olm & 3.90 & \textbf{+45.64\%} & +54.08\% & +15.34\% & +90.77\% \\ 
        ~ & \gmtwo & 5.97 & \textbf{+22.45\%} & +85.37\% & +19.03\% & +94.97\% \\ 
        ~ & \gmsev & 10.91 & +18.97\% & +42.45\% & \textbf{+20.87\%} & +120.99\% \\ 
        ~ & \lmate & 12.99 & \textbf{+28.87\%} & +82.92\% & +17.45\% & +81.60\% \\
        ~ & \lmsev & 18.18 & +16.66\% & +81.18\% & \textbf{+17.91\%} & +83.20\% \\
        ~ & \mis & 9.35 & \textbf{+45.99\%} & +78.86\% & +20.10\% & +67.27\% \\ 
        ~ & \mam & 3.12 & -23.72\% & +83.01\% & \textbf{+18.29\%} & +216.03\% \\ 
        \midrule
        \multirow{7}{*}{\bf AQuA} & \olm & 23.08 & \textbf{+23.35\%} & +53.27\% & +18.45\% & +38.60\% \\ 
        ~ & \gmtwo & 15.38 & +15.99\% & +96.52\% & \textbf{+18.63\%} & +68.79\% \\ 
        ~ & \gmsev & 23.08 & \textbf{+19.67\%} & +37.63\% & +16.62\% & +31.63\% \\ 
        ~ & \lmate & 11.54 & \textbf{+17.68\%} & +258.46\% & +16.84\% & +242.29\% \\
        ~ & \lmsev & 30.77 & +11.25\% & +13.85\% & \textbf{+12.22\%} & +15.15\% \\ 
        ~ & \mis & 23.54 & \textbf{+19.75\%} & +67.14\% & +17.80\% & +51.44\% \\ 
        ~ & \mam & 15.38 & -7.67\% & +95.85\% & \textbf{+10.21\%} & +91.81\% \\ 
        \midrule
        \multirow{7}{*}{\bf GSM8k} & \olm & 25.71 & +12.29\% & +99.04\% & \textbf{+13.22\%} & +96.42\% \\ 
        ~ & \gmtwo & 5.91 & +3.38\% & +64.91\% & \textbf{+20.81\%} & +155.84\% \\ 
        ~ & \gmsev & 25.37 & \textbf{+20.10\%} & +89.27\% & +12.61\% & +79.90\% \\ 
        ~ & \lmate & 37.71 & \textbf{+20.53\%} & +20.03\% & +13.66\% & +21.03\% \\
        ~ & \lmsev & 55.55 & +19.15\% & +35.54\% & \textbf{+21.11\%} & +32.33\% \\
        ~ & \mis & 29.66 & +20.18\% & +87.57\% & \textbf{+20.40\%} & +86.01\% \\ 
        ~ & \mam & 3.12 & +5.45\% & +113.46\% & \textbf{+28.85\%} & +68.59\% \\ 
        \midrule
        \multirow{7}{*}{\bf PIQA} & \olm & 75.89 & +17.35\% & +31.77\% & \textbf{+20.04\%} & +31.77\% \\ 
        ~ & \gmtwo & 77.17 & +16.77\% & +22.46\% & \textbf{+20.41\%} & +22.46\% \\ 
        ~ & \gmsev & 79.89 & +18.02\% & +25.17\% & \textbf{+19.40\%} & +25.17\% \\ 
        ~ & \lmate & 80.43 & +16.26\% & +17.49\% & \textbf{+17.02\%} & +17.49\% \\
        ~ & \lmsev & 80.98 & +12.61\% & +13.67\% & \textbf{+13.67\%} & +13.67\% \\ 
        ~ & \mis & 79.35 & \textbf{+20.82\%} & +26.02\% & +16.52\% & +26.02\% \\ 
        ~ & \mam & 73.91 & +1.10\% & +27.86\% & \textbf{+1.62\%} & +27.86\% \\ 
        \midrule
        ~ & Average $\Delta$ & ~
        & +16.39\% & +64.46\% & \textbf{+17.11\%} & +71.40\% \\ 
        \bottomrule
    \end{tabular}
    \label{tab:reasoning}
\end{table*}

Table \ref{tab:reasoning} shows results when applying consistency improving methods to reasoning tasks. The best (relative) performance gains for a particular task is \textbf{bolded}. We find that average performance gain obtained from ``Probabilistic Tokenization'' is higher than the average performance gain obtained by ``CoT + SC''. Interestingly, the ``Probabilistic Oracle'' is significantly better than ``CoT + SC Oracle''. By manually inspecting the reasoning traces, we find that ``Probabilistic'' tokenization is able to produce significantly diverse reasoning paths while ``CoT + SC'' reasoning paths are often just syntactically different. This can be advantageous as it gives the model more chances to improve the ``Oracle'' accuracy. 

In our experiments, the only non-transformer based model i.e. \mam\ showed mixed results when ``CoT + SC'' was applied. While the model was able to sample diverse reasoning paths, majority of them were often wrong or incomplete. Probabilistic tokenization on the other hand was able to 
consistently outperform ``CoT + SC''.

\subsection{Error Analysis}\label{sec:error_analysis}
\renewcommand{\cellalign}{tc}
\renewcommand{\theadalign}{tc}
\begin{table*}[!ht]
    \centering
    \caption{\textbf{Comparison of reasoning paths.} ``CoT + SC'' paths often have superficial linguistic diversity while ``Probabilistic Tokenization'' paths have meaningful diversity in the reasoning paths.}
    \begin{tabular}{p{0.075\textwidth}p{0.85\textwidth}}
        \toprule
        \scriptsize{Question} & \scriptsize{Sasha notices that prices for lumber have gone up 50\% in the last few months after she bought some lumber. Since she has leftovers, she decides to sell them. She has leftover ten 2 x 4 x 10 boards that cost her \$10 each. She also has five 4 x 4 x 10 boards she bought for \$16 each. If she sells them all, how much profit does she make?} \\
        \scriptsize{Answer} & \scriptsize{90}\\
        \midrule
        \makecell{\scriptsize{CoT + SC}\\\scriptsize{Path 1}} & \scriptsize{Sasha has 10 2 x 4 x 10 boards that cost her \$10 each. So she has 10 x 10 = \$100 worth of 2 x 4 x 10 boards. She also has 5 4 x 4 x 10 boards she bought for \$16 each. So she has 5 x 16 = \$80 worth of 4 x 4 x 10 boards. In total, she has \$100 + \$80 = \$180 worth of lumber. Since prices went up 50\%, she can sell them for 1.5 times as much. 1.5 x 180 = 270. The answer is \$270.}\\
        \makecell{\scriptsize{CoT + SC}\\\scriptsize{Path 2}}& \scriptsize{Sasha has ten 2 x 4 x 10 boards that cost her \$10 each. So she has 10 x \$10 = \$100 worth of 2 x 4 x 10 boards. She also has five 4 x 4 x 10 boards she bought for \$16 each. So she has 5 x \$16 = \$80 worth of 4 x 4 x 10 boards. In total, she has \$100 + \$80 = \$180 worth of lumber. Since prices have gone up 50\%, she can sell them for 1.5 times the price she bought them for. So she can sell them for 1.5 x \$180 = \$270. The answer is \$270.}\\
        \midrule
        \makecell{\scriptsize{Probabilistic}\\\scriptsize{Path 1}}& \scriptsize{Sasha has 10 boards that cost her 10 dollars each. She also has 5 boards that cost her 16 dollars each. So she spent 10 x 10 + 5 x 16 = 100 + 80 = 180. Since prices went up 50\%, she can sell the boards for 1.5 times the price she bought them for. So she can sell them for 1.5 x 180 = 270. She makes 270 - 180 = 90 dollars profit. The answer is 90.} \\
        \makecell{\scriptsize{Probabilistic}\\\scriptsize{Path 2}} & \scriptsize{Sasha bought ten 2 x 4 x 10 boards for \$10 each. So she spent 10 x 10 = \$100. She also bought five 4 x 4 x 10 boards for \$16 each. So she spent 5 x 16 = \$80. In total, she spent \$100 + \$80 = \$180. Since prices went up 50\%, she can sell the boards for 1.5 times the price she bought them for. So she can sell the 2 x 4 x 10 boards for 1.5 x 10 = \$15 each. She can sell the 4 x 4 x 10 boards for 1.5 x 16 = \$24 each. So she can sell the 2 x 4 x 10 boards for 10 x 15 = \$150. She can sell the 4 x 4 x 10 boards for 5 x 24 = \$120. In total, she can sell the boards for \$150 + \$120 = \$270. So she makes \$270 - \$180 = \$90 profit. The answer is \$90.} \\
        \bottomrule
    \end{tabular}
    \label{tab:anecdotes}
\end{table*}

To better understand the behavior of probabilistic tokenization, we study the cases where ``CoT + SC'' and ``Probabilistic Tokenization'' predictions do not match each other.

In Table \ref{tab:anecdotes}, we show a representative example from GSM8k. The model used for generating these reasoning paths is \lmate. We find that the 2 reasoning paths sampled by ``CoT + SC'' are logically equivalent. They use the same steps to in the calculation of total selling price. Importantly, they both stop early after that and get incorrect answer due to stopping early. This is a common failure mode for ``CoT + SC'' in reasoning tasks. As opposed to this, ``Probabilistic Tokenization'' is able to generate logically different reasoning paths. The second path sampled using probabilistic tokenization individually calculates the selling prices for different boards while second path calculates the total selling price directly from cost price. Notably, the difference in prompt tokenization also changes how the currency symbol is being used. In many sequences where the model verbalizes the currency (``dollar'' as opposed to ``\$'') in the reasoning paths, we noticed that the input sequence had tokenized each digit in the value separately i.e. ``\$16'' would be tokenized as ``\$$\rvert$1$\rvert$6'' as opposed to BPE's ``\$$\rvert$16''.

\section{Conclusion}
In this work, we propose ``Probabilistic Tokenization'' as a method to improve self consistency in LLMs. We present a principled way to sample multiple tokenizations of a given string using existing BPE tokenizers of pretrained LLMs. As probabilistic tokenization is an input transformation method rather than an output manipulation method, it can be generally applied to any task. 
We find that probabilistic tokenizations offers significant and consistent gains over baseline in 4 reasoning based tasks. Our analysis shows that the primary reason for success of probabilistic tokenization on reasoning tasks is its ability to generate logically diverse reasoning paths. We hope that probabilistic tokenization can be useful for providing deeper insights into the LLM's reasoning process and true abilities (``Oracle'' choices), paving the way for more robust and interpretable AI systems. 

\section{Limitations}
\label{sec:limitations}
In order to estimate the unigram probabilities $p(t^j)$ from Equation \ref{eq:unigram}, we need access to a sufficiently large and diverse dataset of documents. While we used a sample of 10B tokens from a large scale web corpus, this may not always be the optimal choice for all the tasks. On domain specific tasks such as medical question answering or code generation, a web corpus might not be appropriate. Availability to such corpus is essential since errors in estimating $p(t^j)$ can result in suboptimal tokenizations which can hurt performance \citep{jain2023baseline}. On many of the general purpose tasks, this limitation can be addressed by making use of high quality, open source web corpora such as \textsc{FineWeb} \citep{penedo2024fineweb} or \textsc{Dolma} \citep{dolma}. Even in the cases where the actual corpus is not available but token level counts are available, our method can still be applied.

Furthermore, the model capability analysis (Appendix \ref{app:model_scale}) shows that improvements from probabilistic tokenization are greater for more capable models. If the base model is not very capable, it may not be robust to changes in tokenizations or generate superficially diverse reasoning paths. In such cases, probabilistic tokenization may even \textit{hurt} the performance rather than improving. We highlight that this limitation is also present with other methods using chain-of-thought prompting or self-consistency.




\bibliographystyle{plainnat}
\bibliography{refs}







\appendix

\section{Appendix}

\subsection{Related Work}

We compare probabilistic tokenization to works in multi view learning, tokenization based improvements in language modeling and methods improving self consistency.

\paragraph{Consistency Enhancements using Input Transforms} We use a simple 2-layer neural network as our classifier to select a class given likelihood and selected option from multiple tokenizations. In principle, one can follow the rich literature on ensemble diversity \citep{stickland2020diverse, yeo2021robustness,bo2018coteaching} to build a more sophisticated reranker for both log-likelihood and reasoning based tasks that is aware of the transformed input. \citep{hao2019visual} explores similar ideas to improve consistency of outputs in computer vision.

\paragraph{Tokenization Methods to Improve LLM Performance}
Tokenization plays a crucial role in the reasoning and domain understanding capabilities for LLMs \citep{dagan2024getting, singh2024tokenization}. The popular BPE \citep{sennrich2016neural, gage1994} and Unigram \citep{kudo2018subword} tokenizers are still being studied for better understanding \citep{zouhar2023formal}. Some recent works argue that BPE might not be an optimal tokenization method for all tasks or domains \citep{liu2023task, ali2024tokenizer}. Our work is orthogonal to these directions since we do not aim to modify the existing tokenizer and LLM in any way. 

\paragraph{Self Consistency and Diversity of Thoughts in Reasoning}
Several works improve LLM reasoning capabilities using self consistency and chain-of-thought prompting\citep{wei2022chain,kojima2022large,wang2023selfconsistency, yao2023tree}. Following this, many works improve self consistency by either changing the stopping criteria  \citep{aggarwal2023lets, li2024escape} or by designing a sophisticated voting fucntion to replace majority voting \citep{li2024turnig, jain2024lightweight}. Diversity of thoughts (reasoning paths) is also shown to be an important factor limiting LLM's reasoning ability \citep{li2023making,naik2024diversity}. Our work is relevant in this direction as it aims to improve the diversity in reasoning paths using tokenization.

\subsection{Effect of model scale}\label{app:model_scale}

\begin{table*}[!ht]
    \centering
    \caption{\textbf{Comparing effectiveness of probabilistic tokenization at various model parameter scales.} Average task performance is reported for each method with percent improvement over ``Baseline'' in the bracket.}
    \begin{tabular}{lrrrrr}
        \toprule
        Model & Baseline & CoT + SC & CoT + SC & Probabilistic  & Probabilistic \\ 
         ~ & ~ & Majority & Oracle & Majority & Oracle \\
        \midrule
        \gmtwo\ & 26.11 & 29.93{\tiny(+14.65\%)} & 43.68{\tiny(+67.31\%)} & 31.26{\tiny(+19.72\%)} & 48.43{\tiny(+85.51\%)}\\
        \gmsev\ & 34.81 & 41.49{\tiny(+19.19\%)} & 51.74{\tiny(+42.54\%)} & 40.86{\tiny(+17.38\%)} & 57.24{\tiny(+64.42\%)}\\
        \lmate\ & 35.67 & 43.10{\tiny(+20.83\%)} & 69.45{\tiny(+94.73\%)} & 41.46{\tiny(+16.24\%)} & 67.98{\tiny(+90.60\%)}\\
        \lmsev\ & 46.37 & 53.29{\tiny(+14.92\%)} & 63.09{\tiny(+36.06\%)} & 53.89{\tiny(+16.23\%)} & 63.10{\tiny(+36.09\%)}\\
        \bottomrule
    \end{tabular}
    \label{tab:model_scale}
\end{table*}

We study the effectiveness of probabilistic tokenization at various model parameter scales. As shown in Table \ref{tab:model_scale}, we find that probabilistic tokenization robustly improves the performance across many model parameters scale. We find that relative improvements are highest at 7-8B parameter range, roughly when the ``reasoning'' capabilities start to emerge \citep{brown2020language}. At smaller parameter counts, less gains could be explained by the model's weaker reasoning abilities \citep{brown2020language}. Similarly, larger models that are already capable of generating diverse reasoning paths (as evidenced by lower improvements in ``Oracle'' numbers) show comparable gains over ``Baseline'' with both ``CoT + SC'' and ``Probabilstic'' methods. 

\subsection{Counting vs EM for estimating $p(t^j)$}\label{app:counting_vs_em}
As opposed to using EM to estimate $p(t^j)$ in Equation \ref{eq:unigram}, we resort to a simpler counting based method. We acknowledge that this could result in somewhat distorted unigram probabilities but is significantly faster as it does not have to enumerate over $\xalltok$. We provide additional results (Table \ref{tab:em_vs_counting}) on applying EM to a smaller, 100M token subset to conclude that both methods perform comparably. 

\begin{table*}[!ht] 
    \centering
    \caption{\textbf{Comparing effect of using counting vs EM to estimate EM probabilities.} Average task performance is reported. Both methods perform comparably on a \gmtwo\ model.}
    \begin{tabular}{lrrrrr}
        \toprule
        \multicolumn{6}{c}{Reasoning Tasks}\\
        \midrule
        Model & Baseline & CoT + SC & CoT + SC & Probabilistic  & Probabilistic \\ 
         ~ & ~ & Majority & Oracle & Majority & Oracle \\
        \midrule
        \gmtwo\ (EM) & 26.11 & 29.13 & 38.68 & 29.26 & 45.45\\
        \gmtwo\ (Counting) & 26.11 & 27.93 & 37.68 & 29.81 & 45.45\\
        \midrule
        \multicolumn{6}{c}{Loglikelihood Tasks}\\
        \midrule
        Model & Baseline & Most Likely & Majority & Classifier & Oracle\\ \midrule
        \gmtwo\ (EM) & 39.00 & 38.13 & 41.49 & 49.69 & 54.01\\
        \gmtwo\ (Counting) & 39.00 & 37.63 & 41.58 & 51.01 & 55.56\\
        \bottomrule
    \end{tabular}
    \label{tab:em_vs_counting}
\end{table*}

\subsection{Fixed $l$ vs $l\rightarrow\infty$ for sampling tokenizations}\label{app:finite_vs_infinite}
We compare using fixed vs infinite $l$ for sampling a tokenization in Table \ref{tab:fixed_vs_infinite}. Our findings suggest that considering a fixed window of top-$l$ tokenizations may not offer sufficient diversity leading to redundant generations which explain lesser improvements in ``Oracle'' as well as ``Majority'' numbers.
\begin{table*}[!ht] 
    \centering
    \caption{\textbf{Comparing effect of using counting vs EM to estimate EM probabilities.} Average task performance is reported. Both methods perform comparably on a \gmtwo\ model.}
    \begin{tabular}{lrrrrr}
        \toprule
        \multicolumn{6}{c}{Reasoning Tasks}\\
        \midrule
        Model & Baseline & CoT + SC & CoT + SC & Probabilistic  & Probabilistic \\ 
         ~ & ~ & Majority & Oracle & Majority & Oracle \\
        \midrule
        \gmtwo\ ($l=m^2$) & 26.11 & 27.13 & 33.18 & 26.26 & 35.35\\
        \gmtwo\ ($l\rightarrow\infty$) & 26.11 & \textbf{29.93} & \textbf{43.68} & \textbf{31.26} & \textbf{48.43}\\
        \midrule
        \multicolumn{6}{c}{Loglikelihood Tasks}\\
        \midrule
        Model & Baseline & Most Likely & Majority & Classifier & Oracle\\ \midrule
        \gmtwo\ ($l=m^2$) & 39.00 & \textbf{38.98} & 39.19 & 45.16 & 46.15\\
        \gmtwo\ ($l\rightarrow\infty$)& 39.00 & 38.64 & \textbf{42.98} & \textbf{53.96} & \textbf{56.06}\\
        \bottomrule
    \end{tabular}
    \label{tab:fixed_vs_infinite}
\end{table*}


\subsection{Resources Used}
\label{sec:resources_used}
We used a single NVIDIA A100 GPU with a 64 core AMD CPU to run our inferences. The estimated total GPU hours is 600 hours. Our implementation is based on the \texttt{lm-evaluation-harness}: \url{https://github.com/EleutherAI/lm-evaluation-harness} and \texttt{sentencepiece}: \url{https://github.com/google/sentencepiece}. 
We use 7 models for our experiments, all of which are opensource. The list of model and the URL with checkpoints available and licenses are listed below:
\begin{itemize}
    \item[] \olm\ : \url{https://huggingface.co/allenai/OLMo-7B} {\bf License:} Apache-2.0
    \item[] \gmtwo\ : \url{https://huggingface.co/google/gemma-2b} {\bf License:} Gemma
    \item[] \gmsev\ : \url{https://huggingface.co/google/gemma-7b} {\bf License:} Gemma
    \item[] \lmate\ : \url{https://huggingface.co/meta-llama/Meta-Llama-3-8B} {\bf License:} llama3
    \item[] \lmsev\ : \url{https://huggingface.co/meta-llama/Meta-Llama-3-70B} {\bf License:} llama3
    \item[] \mis\ : \url{https://huggingface.co/mistralai/Mistral-7B-v0.1} {\bf License:} Apache-2.0
    \item[] \mam\ : \url{https://huggingface.co/state-spaces/mamba-2.8b-hf} {\bf License:} Apache-2.0
\end{itemize}

Subsequently we used \textsc{FineWeb} dataset available at \url{https://huggingface.co/datasets/HuggingFaceFW/fineweb} to collect frequencies of the tokens in the for probabilistic tokenization mentioned in Section \ref{sec:probtok}.

\end{document}